OpenStudy (anonymous):

-4x + 5y = -11 2x + 3y = 11 When I solved I got (4,1) Could someone show me how you solve and what you get.

OpenStudy (anonymous):

try to eliminate one of the variables.... what happens if you multiply the bottom equation by 2 and then add it to the top equation?

OpenStudy (phoenixfire):

Solve one equation for a variable. 2x+3y=11 2x=11-3y x=(11-3y)/2 substitute into other equation -4x+5y=-11 -4((11-3y)/2)+5y=-11 ((-44+12y)/2)+5y=-11 -22+6y+5y=-11 11y=11 y=1 Now go back to any of the equations and substitute y=1 and solve for x -4x+5y=-11 -4x+5(1)=-11 -4x=-16 x=16/4 x=4 so your answer is correct at (4, 1)

OpenStudy (anonymous):

lol 2*(2x + 3y = 11) = 4x + 6y = 22 + -4x + 5y = -11 ---------------- 11y =11 y=1 2x+3=11 2x=8 x=4